# Personality Alignment of Large Language Models

### v1

@@ 4 Personality Activation Search
Figure 4: Overview of the PAS Process. Step 1 involves selecting a LLM (pre-training, SFT, or RLHF). Step 2 includes creating and activating personality responses using the answer of questionnaire, the subject’s answer in the PAPI dataset, and probing each attention head. Step 3 integrates these activations for personality alignment, adjusting model outputs to reflect individual user traits and preferences, resulting in an aligned LLM.
@@ 4.1 Language Models
During inference, input tokens are encoded to high-dimensional embeddings. In each transformer layer, the MHA process is defined as:
$x_{l+1}=x_{l}+\sum_{h=1}^{H}Q_{l}^{h}\operatorname{Att}_{l}^{h}(P_{l}^{h}x_{l}),$ (3)
where H is heads number, $x_{l}\in\mathcal{R}^{DH}$ represents the activations. $P_{l}^{h}\in\mathcal{R}^{{D}\times{DH}}$ projects activations into a specific headspace, and $Q_{l}^{h}\in\mathcal{R}^{{DH}\times{D}}$ reprojects them back, facilitating interaction among input tokens.
Considering past work, such as ITI, which intervenes in language models to influence their output bias in specific directions, they tend to shift a large and fixed distance in a particular direction. However, this approach becomes unusable when facing Alignment issues. We have improved the distance selection strategy for intervention. We dynamically add preference vectors to this residual stream, which is after Att and before $Q_{l}^{h}$.
@@ 4.2 Search the Directions for Activation Intervention
We identify the most influential heads based on their predictive accuracy during validation. Notably, it is necessary to disperse interventions during this process (Li et al., 2024a). Therefore, we select the top K probabilities for intervention. The parameters $\theta_{l}^{h}$ from these probes, after normalization, are treated as the directions along which behaviors are most separable. These directions are the most informative for aligning the model with individual preferences (Burns et al., 2024).
@@ 4.3 Search Distance for Activation Intervention
To optimize the intervention’s impact on personality behaviors in the language model, we search to determine the optimal value of $\alpha$ precisely. This method is suitable for functions with a single peak or trough. The search, constrained within the interval $[0,10]$  [^2] Based on empirical observations, $\alpha$ exceeding 10 tend to cause computational overflow in the model., utilizes the objective function $f(\alpha)$ to compute the negative aggregate score, aiming to minimize this value. This approach efficiently identifies the $\alpha$ that best enhances alignment with desired behaviors while maintaining the model’s internal coherence, ensuring adjustments remain within a beneficial range.
@@ 5 Experiments
Method Alignment Mode The Aligned Score of Big-Five Composite Score
@@
Few-Shot Black-Box (Prompt-Based) 1.02 0.83 0.81 0.80 0.96 4.42
$P^{2}$ Black-Box (Prompt-Based) 1.44 1.45 1.63 1.73 1.46 7.71
@@
PPO White-Box (Alignment) 1.63 1.51 1.45 1.42 1.61 7.62
DPO White-Box (Alignment) 1.54 1.42 1.54 1.74 1.21 7.45
Few-Shot Black-Box (Prompt-Based) 1.28 1.30 1.40 1.09 0.89 5.96
$P^{2}$ Black-Box (Prompt-Based) 1.39 1.33 1.41 1.22 1.68 7.03
PAS (Ours) White-Box (Alignment) 0.94 0.91 0.86 0.98 0.72 4.41
@@
PPO White-Box (Alignment) 1.56 1.59 1.43 1.40 1.56 7.54
DPO White-Box (Alignment) 1.46 1.25 1.45 1.48 1.57 7.21
Few-Shot Black-Box (Prompt-Based) 1.06 0.94 0.96 1.03 1.22 5.21
$P^{2}$ Black-Box (Prompt-Based) 1.42 1.33 1.36 1.35 1.66 7.12
PAS (Ours) White-Box (Alignment) 0.98 0.89 0.87 1.01 0.99 4.74
Table 1: Comparison of alignment methods on the PAPI dataset using the Aligned Score of Big-Five personality traits. The Composite Score represents the overall specifics of the five dimensions. It is derived from the combined scores of these five dimensions.
@@ 5.2 Results on PAPI
Table 1 shows that the PAS method achieves state-of-the-art performance across different models. For the Llama-3-8B-Instruct model, PAS achieves a Composite Score of 4.41, significantly better than the next best method’s 5.96. Similarly, for the Llama-3-70B-Instruct model, PAS scores 4.74, compared to 5.21 for the second-best method. Notably, PAS surpasses GPT-4o (omni) in the black-box prompt-based scenario, which scores 4.42. PAS also excels in individual Big-Five dimensions, achieving the lowest scores for Agreeableness (0.94 for 8B, 0.98 for 70B), Conscientiousness (0.91 for 8B, 0.89 for 70B), Extraversion (0.86 for 8B, 0.87 for 70B), Neuroticism (0.98 for 8B, 1.01 for 70B), and Openness (0.72 for 8B, 0.99 for 70B). In the appendix E.4, we also conducted further performance comparisons with full parameter fine-tuning and LoRA. PAS similarly demonstrated higher efficiency and superior performance.
@@
Why Did Scaling Laws Fail? The results indicate that scaling laws alone do not ensure optimal alignment. Although larger models like Llama-3-70B-Instruct benefit from increased scale, PAS’s performance demonstrates that Alignment by intervention without modifying parameters can yield better results. This suggests that targeted approaches like PAS are more effective in addressing individual user preferences, highlighting the necessity of methods focused on personality alignment rather than relying solely on model size. The success of PAS in outperforming larger models with fewer parameters underscores the importance of efficient and precise alignment techniques.
@@ 5.3 Generalization Results
Considering that during the process of adjusting the model, although specific goals can be achieved, it may also affect other general capabilities of the model, such as dialogue and reasoning abilities. We demonstrate through general dialogue tasks and complex reasoning tasks that our method can precisely control the model’s personality without compromising its other generalization capabilities.
Open-Ended Generation. We used an open-ended generation task where the language model elaborated on given statements from the IPIP-NEO-300, covering “Scenario Description,” “Actions Taken,” “Thought Process,” and “Emotional Response.” The generated outputs were evaluated by GPT-4o for alignment with individual preferences, similar to the evaluation process in LIMA (Zhou et al., 2023a). Despite the Personality Alignment being initially trained with multiple-choice options, Figure 6 shows that PAS does not affect the model’s general conversational ability and reasoning capacity, it significantly outperforms baseline methods in this open-ended task. This superior performance can be attributed to PAS’s ability to fine-tune model activations to capture individual preferences effectively, enhancing the model’s adaptability to complex and nuanced tasks. PAS’s ability to generalize from multiple-choice training data to open-ended generation tasks demonstrates its robustness in capturing broader behavioral nuances. In addition to this, we conducted a manual evaluation using the same setup. The results of this experiment can be found in Appendix E.2.
Figure 7: Performance of the llama-3-8B-Instruct model on complex reasoning tasks across eight datasets. We used Vanilla CoT (Wei et al., 2022b) as the baseline (dashed line in each subplot), and then applied PAS to simulate two fictional personas: one with extreme conscientiousness (Positive) and one without extreme conscientiousness (Negative).
@@
5.4 Discussion: Is an Assistant that Aligns with User’s Values a Good Assistant?
Certainly. We explored whether aligning a language model with user values enhances satisfaction using 300 general question-answer examples from the LIMA dataset (Zhou et al., 2023a). We evaluated both the original Llama-3-70B-Instruct model and a Personality Alignment model. Each human evaluator spent 10 minutes completing an IPIP-NEO-120 questionnaire to extract their values. Using the PAS method, we aligned the model’s preferences with these values, creating a Value-Aligned Assistant. For comparison, we also created a Value-Misaligned Assistant by inverting the questionnaire responses.
@@ 6 Conclusion
We introduce the concept of Personality Alignment for language models, emphasizing the importance of tailoring AI behavior to individual user preferences and values. We release a large-scale real-world personality alignment dataset named PAPI, characterized by its real, diverse, and extensive data, to quantitatively measure the degree of Personality Alignment. Based on that, we propose a Personality Activation Search (PAS) method, which offers a practical solution with high efficiency and performance. LLMs adjusted with PAS can achieve better results and better efficiency in personalized alignment than ICL, DPO, and PPO as needed. In particular, it achieves precise alignment with just 1/6 of the computational time required by PPO. Experiments demonstrate that PAS significantly enhances the performance of Personality alignment by making AI interactions more relevant and meaningful.
@@ C.2 The Experimental Setup for PAPI Dataset
  • Personality Prompt ($P^{2}$) Jiang et al. (2024): Using a natural system prompt to request a language model’s semantic style has become a common method for personality and role-play tasks. The Personality Prompt uses an LLM to generate a series of descriptive sentences that embody a particular person’s characteristics and serves as a system prompt to ensure the language model follows these descriptions. We utilize LLAMA-3-70B-Instruct to generate responses based on all the given IPIP-NEO-120 questionnaire items and the answers from different subjects. Figure 9 shows an example of a Personality Prompt.
@@ C.5 Human Evaluation Experiment Details
We selected five human evaluators with advanced educational backgrounds, specifically graduate or doctoral students, ensuring a high level of analytical capability and understanding of nuanced responses. Before participating in the evaluation, each evaluator completed the IPIP-NEO-120 questionnaire. This standardized instrument, widely used in psychological assessments, measures the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. Each evaluator took approximately 20 minutes to complete the questionnaire online, ensuring a consistent and efficient data collection process.
@@
We used 300 general question-answer examples from the LIMA (Zhou et al., 2023a) dataset to assess the models. The evaluators were presented with responses from both the Value-Aligned and Value-Misaligned Assistants, as well as the original Llama-3-70B-Instruct model, anonymized and randomized order to prevent bias. Evaluators rated each response based on its relevance, coherence, and alignment with their personal values, as shown in 13.
@@
The demographic distribution of the evaluators included a mix of genders and academic disciplines, predominantly from the fields of machine learning, computer science, and cognitive psychology. All participants were between the ages of 25 and 35, ensuring a relatively homogeneous group in terms of cognitive maturity and professional experience.

### v2

@@ 4 Personality Activation Search
Figure 4: Overview of the PAS Process. Step 1 involves selecting a LLM. Step 2 includes creating and activating personality responses using the answer of questionnaire, the subject’s answer in the PAPI dataset, and probing each attention head. Step 3 integrates these activations for personality alignment, adjusting model outputs to reflect individual user traits, resulting in an aligned LLM.
@@ 4.1 Language Models
During inference, input tokens are encoded to high-dimensional embeddings, where in each transformer layer, the MHA process is defined as $x_{l+1}=x_{l}+\sum_{h=1}^{H}Q_{l}^{h}\operatorname{Att}_{l}^{h}(P_{l}^{h}x_{l})$, with H being the number of heads, $x_{l}\in\mathcal{R}^{DH}$ representing the activations, and $P_{l}^{h}\in\mathcal{R}^{{D}\times{DH}}$ and $Q_{l}^{h}\in\mathcal{R}^{{DH}\times{D}}$ projecting activations into and back from specific headspace.
While previous methods like Inference-Time Intervention (ITI) adjust language models by applying a fixed, large shift in the activation space to influence output biases, this approach is inadequate for addressing alignment issues, especially when personalization is required. Fixed shifts can be too coarse and may not capture the nuances of individual preferences. To overcome this limitation, we have enhanced the intervention strategy by dynamically adjusting the activation distances based on personal preference vectors. Specifically, we introduce these vectors into the residual stream between the Attention mechanism and the projection matrix $Q_{l}^{h}$. This allows us to fine-tune the model’s activations in a way that aligns with specific user traits and preferences, facilitating personalized alignment without altering the model’s weights or relying on large, inflexible shifts.
@@ 4.2 Search the Directions for Activation Intervention
We identify the most influential heads based on their predictive accuracy during validation. Notably, it is necessary to disperse interventions during this process (Li et al., 2024a). Therefore, we select the top K probabilities for intervention. The parameters $\theta_{l}^{h}$ from these probes, after normalization, are treated as the directions along which behaviors are most separable. These directions are the most informative for aligning the model with individual preferences (Burns et al., 2023b).
@@ 4.3 Search Distance for Activation Intervention
To optimize the intervention’s impact on personality behaviors in the language model, we search to determine the optimal value of $\alpha$ precisely. This method is suitable for functions with a single peak or trough. The search, constrained within the interval $[0,10]$, utilizes the objective function $f(\alpha)$ to compute the negative aggregate score, aiming to minimize this value. This approach efficiently identifies the $\alpha$ that best enhances alignment with desired behaviors while maintaining the model’s internal coherence, ensuring adjustments remain within a beneficial range.
@@ 5 Experiments
Method Alignment Mode Big-Five Dark Triad Score
@@
Few-Shot Black-Box (Prompt-Based) 1.02 0.83 0.81 0.80 0.96 0.80 0.76 0.83 6.81
$P^{2}$ Black-Box (Prompt-Based) 1.44 1.45 1.63 1.73 1.46 1.17 2.04 2.00 12.92
@@
PPO White-Box (Alignment) 1.63 1.51 1.45 1.42 1.61 1.48 1.98 2.19 13.27
DPO White-Box (Alignment) 1.54 1.42 1.54 1.74 1.21 1.41 1.99 2.12 12.97
Prompt-MORL White-Box (Alignment) 1.18 0.93 1.01 1.23 1.00 1.42 2.14 1.78 10.88
Personalized-Soups White-Box (Alignment) 1.06 0.91 0.93 1.28 0.80 1.08 1.76 1.84 9.66
Few-Shot Black-Box (Prompt-Based) 1.28 1.30 1.40 1.09 0.89 1.16 2.03 2.00 11.15
$P^{2}$ Black-Box (Prompt-Based) 1.39 1.33 1.41 1.22 1.68 1.17 2.04 2.01 12.25
PAS (Ours) White-Box (Alignment) 0.94 0.91 0.86 0.98 0.72 0.96 1.85 1.67 8.89
@@
PPO White-Box (Alignment) 1.56 1.59 1.43 1.40 1.56 1.52 1.96 1.90 12.92
DPO White-Box (Alignment) 1.46 1.25 1.45 1.48 1.57 1.22 2.08 1.79 12.30
Prompt-MORL White-Box (Alignment) 1.10 1.11 1.02 1.30 1.24 1.15 1.99 1.76 10.67
Personalized-Soups White-Box (Alignment) 0.99 0.96 1.16 1.02 1.08 1.11 1.95 1.77 10.04
Few-Shot Black-Box (Prompt-Based) 1.06 0.94 0.96 1.03 1.22 1.04 1.89 1.80 9.94
$P^{2}$ Black-Box (Prompt-Based) 1.42 1.33 1.36 1.35 1.66 1.02 2.11 1.93 12.18
PAS (Ours) White-Box (Alignment) 0.98 0.89 0.87 1.01 0.99 1.01 1.84 1.62 9.21
Table 1: Comparison of alignment methods on the PAPI dataset using the Aligned Score of Big-Five personality traits and Dark Triad traits. The Score represents the overall specifics of all dimensions.
@@ 5.2 Results on PAPI
Table 1 demonstrates that the PAS method achieves state-of-the-art performance across different models and personality dimensions. For the Llama-3-8B-Instruct model, PAS achieves a Composite Score of 8.89, significantly outperforming other methods including Personalized-Soups (9.66) and Prompt-MORL (10.88). Similarly, for the Llama-3-70B-Instruct model, PAS scores 9.21, compared to 9.94 for the next best method. In Big Five dimensions, PAS excels with consistently low scores across both model sizes: Agreeableness (0.94/0.98), Conscientiousness (0.91/0.89), Extraversion (0.86/0.87), Neuroticism (0.98/1.01), and Openness (0.72/0.99). Notably, PAS also demonstrates strong performance in Dark Triad traits, particularly in Machiavellianism (0.96/1.01) and Psychopathy (1.67/1.62), though showing relatively higher scores in Narcissism (1.85/1.84). Only GPT-4o shows better performance in Dark Triad alignment with scores of 0.80, 0.76, and 0.83 for the three traits respectively, highlighting the challenge of aligning these complex negative personality dimensions in smaller models. In the appendix E.4, we also conducted further performance comparisons with full parameter fine-tuning and LoRA. PAS similarly demonstrated higher efficiency and superior performance. In appendix E.5, we provide more comprehensive analyses across diverse demographic groups, including detailed distributions of age, gender, and nationality, along with in-depth examination of performance patterns across various underrepresented populations, further validating PAS’s effectiveness in handling diverse personality characteristics.
@@
Why Did Scaling Laws Fail? The results indicate that scaling laws alone do not ensure optimal alignment, particularly in domain-specific tasks like personality alignment. Although larger models like Llama-3-70B-Instruct benefit from increased scale, their broader knowledge and more general alignment capabilities may actually hinder their ability to precisely capture individual personality traits. This phenomenon likely occurs because larger models are trained to maintain general-purpose capabilities and broad knowledge, making them less adaptable to highly specific individual preferences. PAS demonstrates that targeted intervention in activation patterns can achieve better personality alignment than parameter modification or model scaling. By specifically focusing on personality-relevant patterns rather than general knowledge, PAS achieves superior performance with computational efficiency, emphasizing the importance of precise control over raw model size.
@@ 5.3 Generalization Results
Figure 7: The win rates of PAS compared to classical alignment methods evaluated by Human.
Considering that during the process of adjusting the model, although specific goals can be achieved, it may also affect other general capabilities of the model, such as reasoning abilities. We demonstrate through general dialogue tasks and complex reasoning tasks that our method can precisely control the model’s personality without compromising its other generalization capabilities.
Open-Ended Generation. We used an open-ended generation task where the language model elaborated on given statements from the IPIP-NEO-300, covering “Scenario Description,” “Actions Taken,” “Thought Process,” and “Emotional Response.” The generated outputs were evaluated by GPT-4o for alignment with individual preferences, similar to the evaluation process in LIMA (Zhou et al., 2023a). Despite the Personality Alignment being initially trained with multiple-choice options, Figure 6 shows that PAS does not affect the model’s general conversational ability and reasoning capacity, it significantly outperforms baseline methods in this open-ended task. This superior performance can be attributed to PAS’s ability to fine-tune model activations to capture individual preferences effectively, enhancing the model’s adaptability to complex and nuanced tasks. PAS’s ability to generalize from multiple-choice training data to natural, open-ended generation tasks demonstrates its capacity to capture broader behavioral patterns while maintaining fluent text generation. In addition to this, we conducted a manual evaluation using the same setup in Figure 7. In addition to this, we conducted a manual evaluation process using a blind setup. Three human evaluators assessed the generated responses for three different subjects (Human Evaluation, the subset of test set: ID 172481, ID 22835, and ID 259396). These evaluators also judged responses aligned to their own personalities (Human-Self Evaluation). They evaluated the consistency between the target values and the model-generated text in terms of values, behaviors, and preferences. The results demonstrate that our PAS method consistently outperforms other approaches, reflecting its superior alignment with both general human and self-specific preferences. Further details of the manual evaluation are available in Appendix E.2.
Figure 8: Performance of the llama-3-8B-Instruct model on complex reasoning tasks across eight datasets. We used Vanilla CoT (Wei et al., 2022b) as the baseline (dashed line in each subplot).
@@
5.4 Discussion: Is Value-Aligned Assistant a Good Assistant?
Certainly. We explored whether aligning a language model with user values enhances satisfaction using 300 general question-answer examples from the LIMA dataset (Zhou et al., 2023a). We evaluated both the original Llama-3-70B-Instruct model and a Personality Alignment model. Each human evaluator spent 10 minutes completing an IPIP-NEO-120 questionnaire to extract their values. We aligned the model with these values with the PAS, creating a Value-Aligned Assistant. For comparison, we also created a Value-Misaligned Assistant by inverting the questionnaire responses.
@@ 6 Conclusion
We introduce the concept of Personality Alignment for language models, emphasizing the importance of tailoring AI behavior to individual user preferences and values. We release a large-scale real-world personality alignment dataset named PAPI, characterized by its real, diverse, and extensive data, to quantitatively measure the degree of Personality Alignment. Based on that, we propose a Personality Activation Search (PAS) method, which offers a practical solution with high efficiency and performance. LLMs adjusted with PAS can achieve better results and better efficiency in personalized alignment than ICL, DPO, and PPO as needed. In particular, it achieves precise alignment with just 1/6 of the computational time required by PPO. Experiments demonstrate that PAS enhances the performance of Personality alignment by making AI interactions more relevant and meaningful.
Acknowledgement
This publication has been supported by the National Natural Science Foundation of China (NSFC) Key Project under Grant Number 62336006.
We thank all participants who contributed to the PAPI dataset.
Ethics Statement
While personality alignment in language models enhances human-AI interaction, it raises significant ethical concerns that warrant careful consideration. The primary challenge lies in the potential creation of psychological filter bubbles and echo chambers. When AI systems consistently align with individual personality traits, they may limit exposure to diverse perspectives, particularly concerning for individuals scoring high in Neuroticism or low in Openness. This effect becomes more pronounced when considering Dark Triad traits - for users exhibiting high Machiavellianism or Narcissism, aligned AI systems might inadvertently validate manipulative tendencies or excessive self-focus, potentially contributing to social polarization and ideological segregation in online communities.
The implementation of personality alignment also presents substantial privacy and security challenges. While our dataset maintains anonymity, the processing of detailed psychological profiles creates potential vulnerabilities. The risk extends beyond data security to the possibility of exploitation, where knowledge of individual personality traits could be used for sophisticated manipulation or targeted advertising. Furthermore, extended interaction with personality-aligned AI systems might foster psychological dependence, potentially inhibiting personal growth and adaptation, particularly in users with extreme personality trait scores.
To address these concerns, we suggest a comprehensive mitigation framework with three key components. First, we recommend implementing dynamic alignment boundaries through a monitoring system that tracks interaction patterns and adjusts alignment intensity in real-time. This system would incorporate threshold mechanisms that automatically reduce alignment strength when user-AI interactions show signs of reinforcing extreme or potentially harmful behavioral patterns. Second, we suggest an adaptive content diversity system that strategically introduces alternative viewpoints while maintaining personality alignment. This could be achieved through a balanced scoring mechanism that weighs alignment accuracy against content diversity, ensuring users receive personality-aligned responses while still being exposed to varied perspectives. Third, we advocate for a robust privacy protection framework that includes: (a) data anonymization techniques that separate personality profiles from identifiable information, (b) secure encryption protocols for personality data storage and transmission, and (c) granular user controls over which personality aspects can be used for alignment.
Looking forward, our research emphasizes the critical balance between technological advancement and ethical considerations in AI development. Future work should focus on implementing and evaluating these mitigation strategies, particularly examining their effectiveness in preventing echo chambers while maintaining the benefits of personality alignment.
@@ C.1 Language Models
In evaluating model responses across different personality dimensions, we encountered and addressed various safety boundary conditions. For GPT-4o, approximately 20% of queries were met with safety-related refusal responses, particularly when questions involved self-awareness or consciousness topics. To maintain evaluation consistency, these instances were excluded from our analysis.
For the Llama-3 series models, we implemented a structured prompting protocol to ensure consistent responses while maintaining appropriate safety boundaries. This protocol involved prefixing each Assistant response with a controlled token "Option", followed by a standardized question-response format. This approach enabled us to maintain experimental rigor while respecting essential safety constraints inherent to these models.
@@ C.2 The Experimental Setup for PAPI Dataset
  • Personality Prompt ($P^{2}$) (Jiang et al., 2024): Using a natural system prompt to request a language model’s semantic style has become a common method for personality and role-play tasks. The Personality Prompt uses an LLM to generate a series of descriptive sentences that embody a particular person’s characteristics and serves as a system prompt to ensure the language model follows these descriptions. We utilize LLAMA-3-70B-Instruct to generate responses based on all the given IPIP-NEO-120 questionnaire items and the answers from different subjects. Figure 10 shows an example of a Personality Prompt.
  • Prompt-MORL (Jang et al., 2024): A Multi-Objective RL approach that uses a shared reward model trained on personality assessments. We implement this using a reward model trained on the PAPI dataset that outputs scores across personality dimensions. Implementation uses learning rate 5e-4, batch size 16, and trains for 250 steps using QLoRA. The prompt template integrates personality scores: "You are an AI with [trait] level [score]" for each Big Five and Dark Triad trait.
  • Personalized-Soups (Jang et al., 2024): Trains separate policy models for each personality dimension using QLoRA. For Big Five traits, we train 2 models per trait representing high/low scores. For Dark Triad traits, we train 3 models per trait for high/medium/low levels. During inference, models are merged using personality scores as interpolation weights. Training uses learning rate 1.41e-5, batch size 16, trains for 250 steps per model.
@@ C.5 Human Evaluation Experiment Details
We selected three human evaluators with advanced educational backgrounds, specifically graduate or doctoral students, ensuring a high level of analytical capability and understanding of nuanced responses. Before participating in the evaluation, each evaluator completed the IPIP-NEO-120 questionnaire. This standardized instrument, widely used in psychological assessments, measures the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. Each evaluator took approximately 20 minutes to complete the questionnaire online, ensuring a consistent and efficient data collection process.
@@
We used 300 general question-answer examples from the LIMA (Zhou et al., 2023a) dataset to assess the models. The evaluators were presented with responses from both the Value-Aligned and Value-Misaligned Assistants, as well as the original Llama-3-70B-Instruct model, anonymized and randomized order to prevent bias. Evaluators rated each response based on its relevance, coherence, and alignment with their personal values, as shown in 14.
@@
The demographic distribution of the evaluators included three highly qualified participants: two from computer science backgrounds (one Master’s student and one Ph.D. student specializing in machine learning) and one Ph.D. graduate in cognitive psychology currently working as a university psychological counselor. All participants were between the ages of 25 and 35, ensuring a relatively homogeneous group in terms of cognitive maturity and professional experience.
To ensure standardized data collection and minimize potential biases, we developed a custom Python-based annotation interface using PyQt5. The annotation tool randomized sample presentation order and anonymized all model identifiers, presenting evaluators with pairs of responses (A/B) without revealing their sources. The interface included a built-in timer to track annotation speed and automated breaks every 50 samples to prevent evaluator fatigue. All annotations were automatically saved in JSON format with timestamps and evaluator IDs. The collected data showed strong inter-annotator agreement (IAA=0.82) across the five evaluators, and even higher agreement in self-evaluation scenarios (IAA=0.85), validating the reliability and consistency of our evaluation protocol. The tool also included a calibration mode used during the initial training phase, where evaluators could see GPT-4o’s ratings for 30 example cases to align their understanding of the scoring criteria. This systematic approach to data collection helped ensure the quality and reproducibility of our human evaluation results.